# Super-resolving frequency measurement with mode-selective quantum memory

### v1

@@ Abstract
High-precision optical frequency measurement is indispensable to modern science and technology, yet conventional spectroscopic techniques struggle to resolve spectral features narrower than their bandwidth. Here, we introduce a unique platform for super-resolving frequency estimation utilizing a mode-selective atomic Raman quantum memory implemented in warm cesium vapor. By precisely engineering the light-matter interaction, our memory coherently stores the optimal temporal mode with high fidelity and retrieves it on-demand, realizing a mode crosstalk as low as 0.34%. For the task of estimating the separation between two spectral lines, we experimentally measure the mean squared error of the frequency estimate, achieving a sensitivity as small as 1/20 of the linewidth with up to a (34±$\pm 4$)-fold enhancement in precision over direct intensity measurements. This enhanced resolution of frequency measurements, combined with the memory’s capability of on-demand storage, retrieval and mode conversion, paves the way for multifunctional memory-based time-frequency sensors and their integration within quantum networks.
@@ I Introduction
The time-frequency (TF) degree of freedom of light underpins a broad range of applications spanning high-resolution spectroscopy [1], precision timekeeping [2, 3], ultrafast optics [4] and emerging quantum technologies [5, 6, 7, 8]. In spectroscopy, for example, resolving spectral lines is essential for probing the atomic and molecular properties of matter. Frequency measurements are also crucial for quantum metrology and sensing [9, 10, 11], as they largely rely on measuring the transition between energy levels of quantized physical systems. However, the precision of these optical measurements has long been constrained by instrumental limitations. Every spectral feature has an intrinsic linewidth set by the Fourier limit - a lower limit being the inverse of either the measurement time or the signal’s temporal duration. The resolving power of conventional spectrometers is compounded by the Rayleigh criterion [12], which defines the minimal resolvable separation between two spectral lines. As the line separation approaches zero, the uncertainty in their resolution escalates rapidly. This phenomenon, often termed Rayleigh’s curse, poses a critical barrier, particularly when probing weak or photon‐limited signals.
@@
The ultimate precision in resolving spectral features is bounded by the quantum mechanical nature of the optical field [13]. From quantum estimation theory, Tsang, Nair, and Lu have shown that Rayleigh’s criterion is not a fundamental limit, but rather an artifact of the direct intensity measurement (DI) strategy - with coherent measurements an arbitrarily small separation can be resolved with a constant, fine precision [14]. To circumvent the classical resolution limit, implementing a coherent mode filter to select the optimal mode basis before detection, instead of DI measurements, is essential [15, 16, 17, 18, 19, 20]. The practical realization of this advantage, however, relies critically on the accuracy of the mode filtering, as experimental imperfections like mode crosstalk and detector noise substantially diminish the achievable precision [21, 22, 23, 24, 25, 26]. The high-fidelity coherent mode filtering is generally needed for broader quantum information processing and quantum metrology technologies [6, 7, 27, 28, 29]. Looking ahead, the future quantum networks demand sensor nodes with such capability of coherent TF processing, ideally integrated with on-demand buffering for signal synchronization [30, 31], coherent bandwidth and frequency conversion to interface different physical platforms and network channels [32, 33, 34] - all while preserving quantum coherence. These integrated functionalities are essential for building robust, distributed quantum-enhanced sensors [35] capable of dynamically adjusting to varying environmental conditions across a network [36, 37, 38, 39].
@@
In TF super-resolution, quantum pulse gates (QPGs), which use nonlinear waveguides for mode-selective frequency conversion, have been deployed to resolve temporal and spectral separations for ultrafast pulses with hundreds of GHz bandwidth [40, 41]. While QPGs demonstrate programmable mode selectivity [42], they inherently lack on-demand storage and buffering capabilities. In the narrow-band frequency regime, time-inversion interferometry using gradient echo memory has demonstrated sub-Rayleigh resolution for tens of kHz bandwidth pulses [43], but they are limited in the ability to select arbitrary modes and require cumbersome magnetic field gradients and cryogenics. Additionally, there have also been works focusing on frequency super-resolution across tens of GHz bandwidth using dispersion engineering and electro-optic modulators [44, 45]. However, no existing platform, including conventional spectroscopic tools, delivers high‐precision super‐resolution in the MHz to GHz bandwidth, while also providing on‐demand storage, retrieval and user-defined mode selectivity.
@@
In this work, we introduce a new scheme of TF super-resolution via high-fidelity coherent mode filtering in atomic Raman quantum memory. Photonic quantum memories have been widely studied as devices capable of absorbing and re-emitting photonic states on demand, with applications in quantum networks, communication and computing [46, 47, 48, 49, 50, 51, 52]. Implemented in a warm cesium vapor, our Raman memory utilizes a three-level atomic system where a strong control field coherently maps an incoming weak signal field onto a collective atomic spin-wave coherence via stimulated Raman absorption. By tailoring the temporal profile of the Raman control pulses, we demonstrate high-fidelity mode-selectivity up to 99.6% for orthogonal Hermite-Gaussian (HG) temporal modes. With this high-fidelity, user-defined coherent mode filter, our memory-based platform stores the optimal signal mode containing the information of spectral line separation and performs detection on the retrieved mode. We calibrate the measurement apparatus and extract frequency separations via a maximum-likelihood estimation procedure applied to the retrieved mode’s photon statistics. Focusing on sub-linewidth frequency separation under various detected photon budgets (from $2\times 10^{3}$ to $100\times 10^{3}$), our platform consistently outperforms DI methods, achieving significant precision enhancements. Operating in the MHz to GHz bandwidth, our memory-based platform extends the toolbox of TF metrology and provides integrated functionalities encompassing mode filtering, buffering, and shaping. These capabilities are ideally suited for the development of next-generation quantum sensor nodes and their integration into future quantum networks.
@@ II.1 Super-resolving measurement in frequency domain
The task of resolving two closely spaced spectral features can be framed as the estimation of a single parameter: the normalized frequency separation $\epsilon=\Delta\omega/\sigma$, where $\Delta\omega$ denotes the frequency difference between the two spectral lines centered at $\omega_{0}\pm\Delta\omega/2$, and $\sigma$ is the spectral linewidth. We consider the signal to originate from two mutually incoherent emitters of equal intensity, each characterized by a Gaussian spectral amplitude $\psi(\omega)=(2\pi\sigma^{2})^{-1/4}\exp(-\omega^{2}/4\sigma^{2})$. Conventional spectroscopy measures the power spectrum, characterized by a normalized intensity profile given by
@@
The separation $\epsilon$ is then estimated by fitting this model to the measured spectrum. However, as $\epsilon\to 0$, the two Gaussian lines increasingly overlap, making it exceedingly difficult to distinguish them and leading to a large estimation error. For this DI strategy, the amount of information about $\epsilon$ extractable from the measurement outcomes is quantified by the Fisher information (FI) per detected photon
[…]
which vanishes in the limit $\epsilon\to 0$ [14]. The precision, quantified by the variance of an unbiased estimator $\hat{\epsilon}$ is lower-bounded by the Cramér–Rao lower bound (CRLB) $\text{Var}(\hat{\epsilon})\geq 1/[N\mathcal{F}(\epsilon)]$, where $N$ represents the number of detected photons. The estimator variance goes to infinity in the limit $\epsilon\to 0$, which suggests that resolving arbitrarily small separations becomes infeasible with DI measurements.
An optimal measurement, in contrast, employs the coherent temporal mode filtering that project the input signal onto an orthonormal set of HG modes. This task is analogous to the two-point resolution problem in spatial imaging. For a Gaussian point spread function, this mode basis has been shown to saturate the quantum FI - the ultimate limit of precision attainable over all possible measurements for a given input quantum state - in the context of separation estimation [14]. Projecting the input signal onto the HG basis in the frequency domain is equivalent to filtering its temporal profile - the Fourier transform of its spectrum - with the corresponding temporal HG basis. The probability of detection in mode $n$ is
$P(n|\epsilon)=\frac{\epsilon^{2n}}{16^{n}n!}\exp\left(-\frac{\epsilon^{2}}{16}\right).$ (1)
And the FI for this HG measurement basis can be calculated as
@@
$\mathcal{F}_{\text{HG}}(\epsilon)=\sum_{n=0}^{\infty}\frac{1}{P(n|\epsilon)}\left(\frac{\partial P(n|\epsilon)}{\partial\epsilon}\right)^{2}\approx\frac{1}{4}.$ (2)
The constant FI saturates the quantum FI for all $\epsilon$, verifying the optimality of the HG measurement basis and enabling super-resolution.
@@
As the HG measurement is particularly advantageous in small separations, we will focus on the precision and ability to resolve small separations under limited photon budgets in realistic experiments. For small separations, the projection onto the $\text{HG}_{0}$ and $\text{HG}_{1}$ mode contributes most of the FI. According to Eq. 1, the ideal projection probabilities onto HG0 and HG1 modes yield $P(1|\epsilon)/P(0|\epsilon)=\epsilon^{2}/16$. This ratio allows for the construction of a simple raw estimator $\hat{\epsilon}_{\text{raw}}=4\sqrt{N_{1}/N_{0}}$, where $N_{0}$ and $N_{1}$ are the experimentally measured counts of HG0 and HG1 projections, respectively.
In practical implementations, the measured projection probabilities will inevitably deviate from the ideal values predicted by Eq. 1 - a discrepancy that can be interpreted as mode crosstalk. To model this, we account for mode crosstalk between HG0 and HG1 measurements via the coupling matrix
$\bm{M}=\begin{pmatrix}\alpha&1-\beta\\ 1-\alpha&\beta\end{pmatrix}$ (3)
@@
where $\alpha,\beta\in[0,1]$ quantify the mode filtering fidelity. The theoretical probabilities $\bm{p}=(P(0|\epsilon),P(1|\epsilon))^{\top}$ are thus perturbed to $\bm{\tilde{p}}=\bm{Mp}$ in the measurement basis, after which $\bm{\tilde{p}}=(\tilde{P}(0|\epsilon),\tilde{P}(1|\epsilon))^{\top}$ is normalized. Assuming low mode crosstalk $\alpha\approx 1$ and $\beta\approx 1$, we derive the FI for small separations using Eq. 2 as (see the Supplementary Materials for derivations)
@@
It can be seen from Eq. 4 that the FI in the presence of mode crosstalk is critically dependent on the leakage from the HG0 mode into the HG1 mode. If $(\epsilon/4)^{2}\ll 1-\alpha$, the FI decreases significantly from the ideal case. Low-crosstalk mode filtering is therefore crucial for resolving very small separations. In the presence of crosstalk, although $\mathcal{F}(\epsilon)\rightarrow 0$ as $\epsilon\rightarrow 0$, Eq. 4 still gives an advantage in FI over DI methods when the crosstalk is small.
To estimate $\epsilon$ in the presence of crosstalk, we employ maximum likelihood estimation (MLE) using the crosstalk-perturbed probabilities $\tilde{P}(0|\epsilon)$ and $\tilde{P}(1|\epsilon)$. The MLE finds the parameter value that maximizes the likelihood function $\mathcal{L}(\bm{x}|\epsilon)=\prod_{i=1}^{N}\tilde{P}(x_{i}|\epsilon)$ for observing the measured data $\bm{x}=\{x_{1},x_{2},\ldots,x_{N}\}$. The estimation is performed subject to the parameter space constraint $\Theta=\{\epsilon\in\mathbb{R}\mid\epsilon\geq 0\}$. For small separations, the MLE estimators can be written as
$\hat{\epsilon}_{\text{MLE}}=\operatorname*{arg\,max}_{\epsilon}[N_{0}\ln{\tilde{P}(0|\epsilon)}+N_{1}\ln{\tilde{P}(1|\epsilon)}].$ (5)
The MLE is asymptotically unbiased and efficient in the limit $N\to\infty$, implying its variance approaches the CRLB.
In real experiments with finite statistics, the MLE for this separation estimation problem exhibits a non-zero bias, particularly at small separations. This bias comprises contributions from both higher-order asymptotic terms [53] and a distinct positive bias arising from the parameter space constraint. In our model, the separation $\epsilon$ is defined in the non-negative domain, $\Theta$. However, due to shot noise, the experimentally observed normalized counts $\bm{f}=(N_{0}/N,N_{1}/N)^{\top}$ may lie outside the physically valid range predicted by our model $\bm{\tilde{p}}$. When this occurs, the likelihood function $\mathcal{L}(\bm{x}|\epsilon)$ takes its maximum at the boundary of the parameter space constraint (i.e. $\epsilon=0$), leading to a positively skewed distribution of the estimator. This skewing effect results in an unavoidable bias for $\epsilon\to 0$, with the magnitude of the bias determined by the estimator’s distribution. The bias diminishes with higher photon counts $N$, as statistical fluctuations on $N_{0}$ and $N_{1}$ decrease, making the estimator’s distribution at the boundary less probable (see the Supplementary Materials for details).
To assess our measurement apparatus, we quantify the estimation performance by the mean squared error (MSE) of the estimator, $\text{MSE}(\epsilon,N)=\braket{(\hat{\epsilon}-\epsilon)^{2}}$. Importantly, the MSE incorporates both the variance and the squared bias of the estimator, $\text{MSE}(\epsilon,N)=\text{Var}(\epsilon)+b(\epsilon,N)^{2}$, being lower bounded by the CRLB with a bias term [54]
@@
is greater than 1 [22]. This minimal resolvable separation defines a modified Rayleigh criterion for resolving spectral lines under certain detected photon budgets. In this work, we employ MSE and PER as the figures of merit to quantify the performance of our experimental platform.

### v2

@@ Abstract
High-precision optical frequency measurement is indispensable to modern science and technology, yet conventional spectroscopic techniques struggle to resolve sub-linewidth spectral features. We introduce a unique platform for super-resolving frequency estimation utilizing a mode-selective atomic Raman quantum memory implemented in warm cesium vapor. By precisely engineering the light-matter interaction, our memory coherently stores the optimal temporal mode with high fidelity and retrieves it on-demand, realizing a mode crosstalk as low as 0.34%. To estimate the separation between two spectral lines, we experimentally measure the mean squared error of the frequency estimate, achieving a sensitivity of 1/20 of the linewidth with a (34±4)-fold enhancement in precision over direct intensity measurements. This enhanced frequency resolution, combined with the memory’s on-demand storage, retrieval, and mode conversion capabilities, paves the way for multifunctional memory-based time-frequency sensors and their integration within quantum networks.
@@ I Introduction
The time-frequency (TF) degree of freedom of light underpins a broad range of applications spanning high-resolution spectroscopy [1], precision timekeeping [2, 3], ultrafast optics [4] and emerging quantum technologies [5, 6, 7, 8]. In spectroscopy, for example, resolving spectral lines is essential for probing the atomic and molecular properties of matter. Frequency measurements are also crucial for quantum metrology and sensing [9, 10, 11], as they largely rely on measuring the transition between quantized energy levels. However, the precision of these optical measurements has long been constrained by instrumental limitations. Every spectral feature has an intrinsic linewidth set by the Fourier limit - a lower limit being the inverse of the measurement time or the signal’s temporal duration. The resolving power of conventional spectrometers is compounded by the Rayleigh criterion [12], which defines the minimal resolvable separation between two spectral lines. As the separation shrinks, the uncertainty in their resolution escalates rapidly. This phenomenon, often termed Rayleigh’s curse, poses a critical barrier, particularly when probing weak or photon‐limited signals.
@@
The ultimate precision in resolving spectral features is bounded by the quantum mechanical nature of the optical field [13]. From quantum estimation theory, Tsang, Nair, and Lu have shown that Rayleigh’s criterion is not a fundamental limit, but rather an artifact of the direct intensity measurement (DI) strategy - with coherent measurements an arbitrarily small separation can be resolved with a constant, fine precision [14]. To circumvent the classical resolution limit, implementing a coherent mode filter to select the optimal mode basis before detection, instead of DI measurements, is essential [15, 16, 17, 18, 19, 20]. The practical realization of this advantage, however, relies critically on the accuracy of mode filtering, as experimental imperfections like mode crosstalk and detector noise substantially diminish the achievable precision [21, 22, 23, 24, 25, 26]. The high-fidelity coherent mode filtering is generally needed for broader quantum information processing and quantum metrology technologies [6, 7, 27, 28, 29]. Looking ahead, the future quantum networks demand sensor nodes with such capability of coherent TF processing, ideally integrated with on-demand buffering for signal synchronization [30, 31], coherent bandwidth and frequency conversion to interface different physical platforms and network channels [32, 33, 34] - all while preserving quantum coherence. These integrated functionalities are essential for building robust, distributed quantum-enhanced sensors [35] capable of dynamically adjusting to varying environmental conditions across a network [36, 37, 38, 39].
@@
In TF super-resolution, quantum pulse gates (QPGs), which use nonlinear waveguides for mode-selective frequency conversion, have been deployed to resolve temporal and spectral separations for ultrafast pulses with hundreds of GHz bandwidth [40, 41]. While QPGs demonstrate programmable mode selectivity [42], they inherently lack on-demand storage and buffering capabilities. In the narrow-band frequency regime, time-inversion interferometry using gradient echo memory has demonstrated sub-Rayleigh resolution for tens of kHz bandwidth pulses [43], but they are limited in the ability to select arbitrary modes and require cumbersome magnetic field gradients and cryogenics. Additionally, there have also been works focusing on frequency super-resolution across tens of GHz bandwidth using dispersion engineering and electro-optic modulators [44, 45]. However, no existing platform delivers high‐precision super‐resolution in the MHz to GHz bandwidth together with on‐demand storage, retrieval and user-defined mode selectivity.
@@
In this work, we introduce a unique scheme of TF super-resolution via high-fidelity coherent mode filtering in atomic Raman quantum memory. Photonic quantum memories have been widely studied for absorbing and re-emitting photonic states on demand, with applications in quantum networks, communication and computing [46, 47, 48, 49, 50, 51, 52]. Implemented in a warm cesium vapor, our platform utilizes a three-level atomic system where a strong control field with tailored temporal profile coherently maps an incoming signal field onto a collective atomic spin-wave coherence via stimulated Raman absorption, achieving high-fidelity mode-selectivity up to 99.6% for orthogonal Hermite-Gaussian (HG) temporal modes. Leveraging this user-defined coherent mode filter, we store the optimal signal mode containing the information of spectral line separation, retrieve it on demand, and apply maximum-likelihood estimation to the retrieved photon statistics to extract frequency separations. Focusing on sub-linewidth frequency separation under various detected photon budgets (from $2\times 10^{3}$ to $1\times 10^{5}$), our platform consistently outperforms DI methods, achieving high precision enhancements. Operating in the MHz to GHz bandwidth, our memory-based platform extends the toolbox of TF metrology and provides integrated functionalities encompassing mode filtering, buffering, and shaping. These capabilities are ideally suited for next-generation quantum sensor nodes and their deployment in quantum networks.
@@ II.1 Super-resolving measurement in frequency domain
We frame the task of resolving two closely spaced spectral features as an estimation of their normalized frequency separation, $\epsilon=\Delta\omega/\sigma$. Here, $\Delta\omega$ denotes the frequency difference between the two spectral lines centered at $\omega_{0}\pm\Delta\omega/2$, and $\sigma$ is the spectral linewidth. We model the sources as two mutually incoherent emitters of equal intensity, each described by a Gaussian spectral amplitude $\psi(\omega)=(2\pi\sigma^{2})^{-1/4}\exp(-\omega^{2}/4\sigma^{2})$. Conventional spectroscopy measures the power spectrum,
@@
The separation $\epsilon$ is then estimated by fitting this model to the measured spectrum. The amount of information about $\epsilon$ extractable from the measurement outcomes is quantified by the Fisher information (FI) per detected photon
@@
$\mathcal{F}_{\text{DI}}(\epsilon)=\int_{-\infty}^{\infty}d\omega\frac{1}{S\left(\omega|\epsilon\right)}\left(\frac{\partial S\left(\omega|\epsilon\right)}{\partial\epsilon}\right)^{2},$
which vanishes in the limit $\epsilon\to 0$ [14]. The precision, quantified by the variance of an unbiased estimator $\hat{\epsilon}$, is lower-bounded by the Cramér–Rao lower bound (CRLB) $\text{Var}(\hat{\epsilon})\geq 1/[N\mathcal{F}(\epsilon)]$, where $N$ represents the number of detected photons. As $\epsilon\to 0$, the two Gaussian lines increasingly overlap and the estimator variance diverges to infinity, making the estimation of arbitrarily small separations infeasible.
@@
An optimal measurement, in contrast, projects the signal onto an orthonormal Hermite–Gaussian (HG) mode basis with coherent filtering. Experimentally, projecting the input signal onto the HG basis in the frequency domain is equivalent to filtering its temporal profile - the Fourier transform of its spectrum - with the corresponding temporal HG basis. This mode basis exhibits a constant FI of $\mathcal{F}_{\text{HG}}(\epsilon)\approx 1/4$ that saturates the quantum FI - the ultimate limit of precision attainable over all possible measurements for a given input quantum state [14]. As the HG measurement is particularly advantageous for small separations, we will focus on the precision and ability to resolve small separations under limited photon budgets in realistic experiments. For small separations, the projection onto the $\text{HG}_{0}$ and $\text{HG}_{1}$ mode contributes most of the FI. The ideal projection probabilities onto HG0 and HG1 modes defines a raw estimator $\hat{\epsilon}_{\text{raw}}=4\sqrt{N_{1}/N_{0}}$, where $N_{0}$ and $N_{1}$ are the experimentally measured counts of HG0 and HG1 projections.
In practical implementations, mode crosstalk modifies the ideal projection probabilities. We can model this effect between the first two HG modes by describing how the ideal probability vector, $\bm{p}$, is perturbed into the measured probability vector, $\bm{\tilde{p}}$. This transformation is defined by a coupling matrix $\bm{M}$, as
$\bm{\tilde{p}}=\bm{Mp}=\begin{pmatrix}\alpha&1-\beta\\ 1-\alpha&\beta\end{pmatrix}\begin{pmatrix}P(0|\epsilon)\\ P(1|\epsilon)\end{pmatrix}$, (1)
@@
where $\alpha,\beta\in[0,1]$ quantify the mode filtering fidelity. The resulting perturbed probability $\bm{\tilde{p}}=\begin{pmatrix}\tilde{P}(0|\epsilon)\\ \tilde{P}(1|\epsilon)\end{pmatrix}$ is then normalized. Assuming low mode crosstalk $\alpha\approx 1$ and $\beta\approx 1$, we derive the FI for small separations as (see Supplementary Information for details)
@@
The FI in the presence of mode crosstalk is highly sensitive to the leakage from the HG0 mode into the HG1 mode, $1-\alpha$. If $(\epsilon/4)^{2}\ll 1-\alpha$, the FI decreases significantly from the ideal case, highlighting the importance of low-crosstalk mode filtering in resolving very small separations.
To estimate $\epsilon$ in the presence of crosstalk, we employ maximum likelihood estimation (MLE) using the perturbed projection probabilities onto the first two HG modes, $\tilde{P}(0|\epsilon)$ and $\tilde{P}(1|\epsilon)$. The MLE finds the parameter value that maximizes the likelihood function $\mathcal{L}(\bm{x}|\epsilon)=\prod_{i=1}^{N}\tilde{P}(x_{i}|\epsilon)$ for observing the measured data $\bm{x}=\{x_{1},x_{2},\ldots,x_{N}\}$. For small separations, the MLE estimator can be written as
$\hat{\epsilon}_{\text{MLE}}=\underset{\epsilon}{\mathrm{argmax}}[N_{0}\ln{\tilde{P}(0|\epsilon)}+N_{1}\ln{\tilde{P}(1|\epsilon)}],$ (3)
subject to the parameter space constraint $\Theta=\{\epsilon\in\mathbb{R}\mid\epsilon\geq 0\}$. The MLE is asymptotically unbiased and efficient in the limit $N\to\infty$, implying its variance approaches the CRLB.
In real experiments with finite statistics, the MLE exhibits a non-zero bias, particularly at small separations. This bias arises from both higher-order asymptotic terms [53] and the non-negativity parameter space constraint $\Theta$. Due to shot noise, the experimentally observed normalized counts $\bm{f}=\begin{pmatrix}N_{0}/N\\ N_{1}/N\end{pmatrix}$ may lie outside the physically valid range. When this occurs, the likelihood function $\mathcal{L}(\bm{x}|\epsilon)$ takes its maximum at the boundary of the parameter space constraint (i.e. $\epsilon=0$), leading to a positively skewed distribution of the estimator. This skewing effect results in an unavoidable bias for $\epsilon\to 0$, with the magnitude of the bias determined by the estimator’s distribution. The bias diminishes with higher photon counts $N$, as statistical fluctuations on $N_{0}$ and $N_{1}$ decrease, making the estimator’s distribution at the boundary less probable (see Supplementary Information for simulation results).
We assess our measurement performance using the mean squared error (MSE), $\text{MSE}(\epsilon,N)=\langle(\hat{\epsilon}-\epsilon)^{2}\rangle$, incorporating both the variance and the squared bias of the estimator, $\text{MSE}(\epsilon,N)=\text{Var}(\epsilon)+b(\epsilon,N)^{2}$. The MSE is lower bounded by the CRLB with a bias term [54]
@@
is greater than 1 [22]. This minimal resolvable separation defines a modified Rayleigh criterion for resolving spectral lines under certain detected photon budgets. In this work, MSE and PER serve as the figures of merit to quantify the performance of our experimental platform.